# !TeX program = pdflatex

### v1

@@ Abstract
Keywords:
large language models natural language processing sentiment analysis marketing electronic word-of-mouth
@@ 1 Introduction
Nowadays, as consumers spontaneously post their opinions regarding products and services on online platforms, such as social media and review apps, user-generated contents (UGCs) has become essential resource for market research in terms of product development, service improvement, and other activities. For example, textual data collected from online platforms and social networks have been utilized for company’s various decision-making, such as product evaluation, feature extraction, and recommendation systems [20, 24, 41, 31].
To extract, utilize, and understand consumer preferences from textual data, pre-processing the data through assigning labels, classifying the text, and evaluating the sentiment is essential. However, these are labor-intensive tasks for humans, whereas automated analysis using natural language processing (NLP) technology requires domain knowledge such as text mining and machine learning. Companies can choose another option to use crowdsourcing such as Amazon Mechanical Turk (MTurk), which has widely used in recent years but is costly, particularly when dealing with large-scale data. Additionally, the quality of crowdsourced data remains a serious concern in academia [3, 12]. Thus, as the enormous amount of data has been accumulated, handling big data becomes more challenging and often impractical.
With the advance of large language models (LLMs), several studies have proposed automated annotation models using LLMs [9, 29, 15]. Compared to human annotators, LLMs have highlighted advantages in processing speed, cost efficiency, and applicability across the wide range of tasks. However, some studies [29] have pointed out the instability of LLM outputs, which can lead to inconsistencies in sentiment predictions. Instead of perceiving this as a limitation, we propose the utilization of this variability through an ensemble approach. By aggregating multiple inferences, the model can improve both stability and accuracy.
Therefore, in this study, we propose a sentiment analysis model that employs ensemble techniques, which not only stabilizes the classification but also reduces the effect of outlier generation of LLM inferences. The remainder of this paper is organized as follows. In Section 2, we review related works to contextualize this study. In Section 3, we introduce our proposed model, followed by an overview and results of the analysis in Section 4. Finally, we discuss the implications and challenges of this study in Section 5.
@@ 2 Related Work
Despite the increasing use of LLMs, few studies have explored ensemble techniques to stabilize predictions. Most prior works have focused on single-run LLM annotations. While a few studies [37, 17] have applied ensemble-like techniques to LLM inference by assigning different characteristics to each LLM, they often rely on hierarchical models that vertically stacks LLMs, where multiple inferences are fed into another LLM for final decision-making. This approach, however, further increases the black-box nature of the model, which makes it difficult to interpret how individual predictions contribute to the final output. Consequently, little research has systematically analyzed the effect of considering the repeated inferences in LLM-based sentiment analysis.
@@ 3.1 Pre-trained Models
We employ Meta AI’s Llama 3 [13] for the pre-trained model, which is trained with more than 15 trillion tokens from publicly available data sources. Furthermore, the fine-tuning [36] process of the model includes more than 10,000 manually annotated data in addition to instructional data, as well as reinforcement learning from human feedback (RLHF) [4]. Considering these learning processes, sufficient accuracy is expected for sentiment analysis of eWOM without additional training.
To reduce the model size, this study employs quantization on pre-trained models and examine the effect of the model scale (i.e., number of parameters in the model). Table 1 lists the employed models. We adopt 70 billion (70B) and 8 billion (8B) models for Llama 3 and 7 billion (7B) model for Llama 2 and quantize each model in 4-bit with GGUF format [14].
@@
  • The‘model id’ represents the corresponding model identifier on Hugging Face (https://huggingface.co/models).
@@ 3.3 Ensemble Strategy
To implement the ensemble learning by repeating the inferences, we focus on the 8B model since 70B model is too costly to be executed. While ordinary annotation process employs majority voting, LLMs may occasionally produce unexpected values such as outliers or missing values. Therefore, as a simple averaging-based method, we adopt the median value across the valid outputs for the aggregation. For sample $i$, the unified output, denoted as $s^{*}_{i}$, is obtained as follows:
@@ 4.1 Dataset Overview
We use Yelp Open Dataset [38], an open dataset publicly available for academic research. Yelp is an online platform on which users post evaluations and reviews about various facilities, including restaurants, stores, and public institutions. A key challenge in analyzing Yelp Dataset is the diversity of restaurants which vary in price range, cuisine, and service. While many studies [1, 2, 22, 26] have analyzed user reviews using this dataset due to its reproducibility, few have employed LLMs for such analyses.
The dataset includes the ratings $rating_{jk}\in\{1,2,\cdots,5\}$ and review texts $review_{jk}$, which are posted by user $j$ about restaurant $k$, where $j\in\{1,2,\dots,J\}$, $k\in\{1,2,\dots,K\}$. A single review was randomly selected for each user. If a user posted a review for the same restaurant multiple times, only the most recent review was considered for sampling. Thus, the sample size $n$ equals the number of users $J$.
@@
Table 2: Summary Statistics of the Data
@@ 4.2 Evaluation
We evaluate model performance using three metrics: concordance rate (Acc.), Pearson’s correlation coefficient (Corr.), and root mean square error (RMSE). Since the target variable is ordinal, we assess not only prediction concordance but also the magnitude of errors using RMSE.
For comparison, we employ six reference models. Among these, we implement three DNN-based models: BERT, Bi-LSTM, and CNN, based on previous studies [6, 23, 34]. The BERT model employs bert-large-uncased for both embedding and classification, while the Bi-LSTM and CNN models incorporate word embeddings (word2vec[25], FastText[19]) and respective architectures. For rule-based methods, we employ VADER, which computes polarity scores directly without training. For machine learning, we adopt a TF-IDF and Linear SVM combination based on previous studies, setting the TF-IDF dimensionality to 4,000. Additionally, we establish a chance-level benchmark by estimating sentiment label distributions from the training data and generating test predictions accordingly.
@@ 4.3 Results and Discussion
The results are shown in Table 3. First, across the non-ensemble models (Model 2–10), most of the LLMs (Model 2–3) except Llama-2 (Model 4) outperformed the reference models in all the metrics. Notably, 70B model (Model 2) demonstrated the high performance across the non-ensemble models. Therefore, among the non-ensemble LLMs, the number of parameters significantly contributes to the accuracy, which is consistent with the previous studies [27]. However, 70B model (Model 2) requires approximately 11 times the processing time of the 8B model (Model 3) for only a 3% increase in accuracy, which means that, even when analyzing only 1,000 samples, a total processing time differs by more than 16 hours. Moreover, all Llama 3 models (Models 2–3) outperformed the Llama 2 model (Model 4). Notably, Llama 2 required more inference time than 80B-parameter Llama 3 model (Model 3), despite having fewer parameters. These results indicate that the model architecture and pre-training process have a significant impact on both the accuracy and processing speed.
@@
Model Name Ensemble Corr. RMSE Acc. Time ($s$)
@@
1. Llama-3-8B-Instruct voting 0.998 0.424 0.778 27.215
2. Llama-3-70B-Instruct none 0.929 0.521 0.779 64.879
3. Llama-3-8B-Instruct none 0.909 0.562 0.749 5.443
4. Llama-2-7B-chat none 0.791 0.860 0.721 5.986
@@
5. BERT (Large) [6] none 0.792 0.941 0.639 -
6. Bi-LSTM [23] none 0.789 0.930 0.636 -
7. CNN [34] none 0.700 1.098 0.596 -
8. VADER [18] none 0.667 1.111    - -
9. Linear SVM [5] none 0.641 1.134 0.627 -
10. Random none 0.019 1.850 0.358 -
@@
from the best model (70B) 7.4% 18.6% -0.1% 58.1%
from the base model (8B) 9.8% 32.7% 3.9% -500.0%
from the best reference 26.0% 54.4% 17.9% -
  • Note. Bold text indicates that the model performs better than all references, while shading behind the index represents the highest performance. Time represents the average duration to process one review.
In such a situation, the proposed ensemble model outperformed the 70B model across correlation and RMSE and almost equivalent in accuracy. This indicates that while the ensemble model achieved similar accuracy to the 70B model, it significantly reduced the margin of error in misclassification with less than half the processing time. Thus, a medium-sized model using ensemble approach with iterative inferences is significantly more efficient in terms of both training time and prediction accuracy than the larger model with a single inference.
Regarding the reference models, although the DNN-based models (Models 5–7) achieved high accuracy as expected, none of them surpassed the performance of Llama 3. Additionally, VADER (Model 8) outperformed the machine-learning model (Model 9). Nevertheless, all models performed above the chance level (Model 10).
@@ 5 Conclusion
From a practical perspective, our findings provide two key implications. First, the ensemble of multiple inferences with a medium-sized model are much faster and more accurate than a single inference with a larger model. This efficiency can enhance the understanding of consumer preference and support quick decision-making for business. Second, since ensemble inference does not require additional training costs, practitioners can readily improve the tasks with this approach in their computational environments.
From an academic perspective, our study suggests that simple ensemble methods can enhance the model performance using the randomness of LLM outputs. Unlike the other ensemble methods, which rely on multiple independently trained models (e.g., bagging or boosting), our approach is based on the stochastic nature of LLM inferences. This indicates that this strategy is generally applicable to other pre-trained LLMs with different architecture or different number of parameters.
Finally, this study presents several challenges for further investigation. First, all the base models employed the same example in the prompt; however, it has been pointed out that the performance of LLMs varies depending on the instructions. Since different examples ensure the more diversity in each model, future research should explore the effects of different prompt configurations across the base models. Second, while this study focus on ensemble of pre-trained models with the same number of parameters, it is necessary to explore the effects of combining pre-trained models of different scales. This study treated all the outputs equally by using the median; however, to effectively integrate models with varying parameter counts, alternative ensemble techniques such as weighted averaging may be required. Third, although this study applied the ensemble of LLM inferences to sentiment analysis on restaurant reviews, the strategy itself is applicable to general text classification tasks using LLMs. Future work should evaluate its effectiveness on broader benchmark across different domains to further validate its generalizability.
Acknowledgment
We are grateful to Yelp for providing access to the dataset. This work was supported by JSPS KAKENHI (Grant No. 24K16472).

### v2

@@ Abstract
Keywords large language models  $\cdot$ natural language processing  $\cdot$ sentiment analysis  $\cdot$ marketing  $\cdot$ electronic word-of-mouth
@@ 1 Introduction
***jniimi@meijo-u.ac.jp
This manuscript has been accepted for the 30th International Conference on Natural Language & Information Systems (NLDB 2025) and will appear in Springer Lecture Notes in Computer Science (LNCS). Codes are available at GitHub: https://github.com/jniimi/ensemble_inference
In the field of marketing, accurate comprehension of customer’s loyalty and preference is crucial as a customer relationship management (CRM) [1, 2]. In particular, as consumers increasingly post opinions on social media and review platform, user-generated contents (UGCs) has become essential resource for market research [3]. Textual data have been utilized for company’s various decision-making, such as product evaluation, feature extraction, and recommendation systems [4, 5, 6, 7, 8].
To extract, utilize, and understand consumer preferences from textual data, pre-processing through assigning labels is essential; however, these are labor-intensive tasks for humans. Manual labelling such as Amazon Mechanical Turk (MTurk) is costly, while traditional natural language processing (NLP) methods require specialized skills. In addition, data quality of crowdsourcing remains a serious concern [9, 10]. Thus, handling big data becomes more challenging and often impractical despite the large amount of accumulated data.
With the advance of large language models (LLMs), several studies have proposed automated annotation models using LLMs [11, 12, 13]. Compared to human annotators, LLMs have highlighted advantages in processing speed, cost efficiency, and applicability across the wide range of tasks. However, the instability of LLM outputs has also been pointed out [13], which can lead to inconsistency of predictions. We therefore facilitate this variability through an ensemble approach by aggregating multiple inferences to acquire both stability and accuracy.
Therefore, this study proposes a sentiment analysis model that employs ensemble techniques, which not only stabilizes the classification but also reduces the effect of outlier generation in LLMs. The remainder of this paper is organized as follows. In Section 2, we review related works to contextualize this study. In Section 3, we introduce our proposed model, followed by an overview and results of the analysis in Section 4. Finally, we discuss the implications and challenges of this study in Section 5.
@@ 2 Related Work
Despite the increasing use of LLMs, few studies have explored ensemble techniques to stabilize predictions. Most prior works have focused on single-run LLM annotations [27]. While a few studies [28, 29] have applied ensemble-like techniques to LLM inference by assigning different characteristics to each LLM, they often rely on hierarchical models that vertically stacks LLMs, where multiple inferences are fed into another LLM for final decision-making. This approach, however, further increases the black-box nature of the model, which makes it difficult to interpret how individual predictions contribute to the final output. Consequently, little research has systematically analyzed the effect of considering the repeated inferences in LLM-based sentiment analysis.
@@ 3.1 Pre-trained Models
Table 1 summarizes the models in this study: Meta AI’s Llama family [30]. We adopt 70 billion (70B) and 8 billion (8B) models for Llama 3 and 7 billion (7B) model for Llama 2. Llama is trained with more than 15 trillion tokens from publicly available data sources. Furthermore, Meta’s instruction-tuning [31] process of the model includes more than 10,000 manually annotated data in addition to instructional data, as well as reinforcement learning from human feedback (RLHF) [32]. Considering these learning processes, sufficient accuracy is expected for sentiment analysis of eWOM. Thus, no additional fine-tuning is performed in this study.
@@
  • The model id represents the corresponding model identifier on Hugging Face (https://huggingface.co/models).
We limit our investigation to the Llama family. Even comparing within the Llama family, we can empirically investigate the effect of the model configuration, such as the model scale (i.e., number of parameters in the model; 7–70B) and version (version 2 and 3).To reduce the computational burden, this study quantizes each model in 4-bit with GGUF format [33] which is widely adopted in LLM community.
@@ 3.3 Ensemble Strategy
To implement the ensemble learning by repeating the inferences, we develop the ensemble extension of 8B model and compare the performance with non-ensemble 70B model. Since the target variable is ordinal scale (1–5 stars), we adopt the median for the ensemble, which is robust to outlier predictions. Majority voting, on the other hand, is commonly employed in ensemble learning; however, it is vulnerable to the case when the number of votes is tied (e.g., two votes for 1, two for 2, and one for 3). Given our focus on the diversity of LLM outputs, the median is more appropriate for the aggregation.
For sample $i\in\{1,2,\cdots,n\}$, the unified output, denoted as $s^{*}_{i}$, is obtained as follows:
@@ 4.1 Dataset Overview
We use Yelp Open Dataset [34], an open dataset publicly available for academic research. Yelp is an online platform on which users post evaluations and reviews about various facilities, including restaurants, stores, and public institutions. Since our proposed method ensembles one-shot LLMs, performance is highly sensitive to pre-training data leakage. The Yelp dataset can only be obtained through an application process, so it is unlikely that an entire corpus was included for pre-training process of Llama family.
A key challenge in analyzing Yelp Dataset is the diversity of restaurants which vary in price range, cuisine, and service. While many studies [35, 36, 37, 38] have analyzed user reviews using this dataset due to its reproducibility, few have employed LLMs for such analyses.
The dataset includes the ratings $rating_{jk}\in\{1,2,\cdots,5\}$ and review texts $review_{jk}$, which are posted by user $j$ about restaurant $k$, where $j\in\{1,2,\dots,J\}$, $k\in\{1,2,\dots,K\}$. A single review was randomly selected for each user. If a user posted a review for the same restaurant multiple times, only the most recent review was considered for sampling. Thus, the sample size $n$ equals the number of users $J$. Since all users and venues are unique, sample size $n=J\times K$.
@@
Table 2: Summary Statistics of the Dataset
@@ 4.2 Evaluation
We evaluate model performance using two metrics: concordance rate (Acc.) and root mean square error (RMSE). Since the target variable is ordinal, we assess not only prediction concordance but also the magnitude of errors using RMSE.
For the model comparison, we employ six reference models for the baseline. Among these, we implement three DNN-based models: BERT, Bi-LSTM, and CNN, based on previous studies [40, 41, 42]. The BERT model employs bert-large-uncased for both embedding and classification, while the Bi-LSTM and CNN models incorporate word embeddings (word2vec[43], FastText[44]) and respective architectures. For rule-based methods, we employ VADER, which computes polarity scores directly without training. For machine learning, we adopt a TF-IDF and Linear SVM combination based on previous studies, setting the TF-IDF dimensionality to 4,000. Additionally, we establish a chance-level benchmark by estimating sentiment label distributions from the training data and generating test predictions accordingly.
@@ 4.3 Results and Discussion
Model Name Ensemble RMSE Acc. Time ($s$)
@@
1. Llama-3-8B-Instruct adopted 0.424 0.778 27.215
2. Llama-3-70B-Instruct - 0.521 0.779 64.879
3. Llama-3-8B-Instruct - 0.562 0.749 5.443
4. Llama-2-7B-chat - 0.860 0.721 5.986
@@
5. BERT (Large) [40] - 0.941 0.639 -
6. Bi-LSTM [41] - 0.930 0.636 -
7. CNN [42] - 1.098 0.596 -
8. VADER [14] - 1.111      - -
9. Linear SVM [46] - 1.134 0.627 -
10. Random - 1.850 0.358 -
@@
from the best single LLM (70B) 18.6% -0.1% 58.1%
from the LLM baseline (8B) 32.7% 3.9% -500.0%
from the best reference model 54.4% 17.9% -
  • Note. Bold metric indicates that the model performs better than all references while cell shading represents the highest performance. Time represents the average duration to process one review.
The results are shown in Table 3. First, non-ensemble LLMs (Model 2–4) outperformed all reference models. Notably, 70B model (Model 2) demonstrated the high performance compared to fewer-parameter models. Therefore, among the non-ensemble LLMs, the number of parameters significantly contributes on performance, which is consistent with the previous studies [45]. However, 70B model (Model 2) requires approximately 11 times the processing time of the 8B model (Model 3) for only a 3% increase in accuracy, which means that, even when analyzing only 1,000 samples, a total processing time differs by more than 16 hours. Moreover, all Llama 3 models (Models 2–3) outperformed the Llama 2 model (Model 4). Notably, Llama 2 required more inference time than 80B model (Model 3), despite having fewer parameters. These results indicate that the model architecture and pre-training process have a significant impact on both the accuracy and processing speed.
In such a situation, the proposed ensemble model even outperformed the 70B model across RMSE and almost equivalent in accuracy. This indicates that, while the ensemble model achieved similar accuracy to the 70B model, it significantly reduced the margin of error in misclassification with less than half the processing time. Thus, a medium-sized model using ensemble approach with iterative inferences is significantly more efficient in terms of both training time and model performance than the larger model with a single inference.
Across the non-LLM baselines (Models 5–10), although the DNN-based models (Models 5–7) achieved high accuracy as expected, none of them surpassed LLM-based models. Additionally, VADER (Model 8) outperformed the machine-learning model (Model 9). Nevertheless, all models performed above the chance level (Model 10).
@@ 5 Conclusion
This study has both academic and practical implications. First, from an academic perspective, our study suggests that simple ensemble methods can enhance the model performance using the randomness of LLM outputs. Unlike the other ensemble methods which rely on multiple independently trained models (e.g., bagging, boosting), our approach is based on the stochastic nature of LLM inferences. This indicates that this strategy is generally applicable to other pre-trained LLMs with different architecture or different number of parameters. Also, from a practical perspective, our findings provide two key implications. First, the ensemble of multiple inferences with a medium-sized model are much faster and more accurate than a single inference with a larger model. This efficiency can enhance the understanding of consumer preference and support quick decision-making for business. Second, since ensemble inference does not require additional training costs, practitioners can readily improve the tasks with this approach in their business environments.
Finally, this study remains several future challenges. First, all the base models employed the same example in the prompt; however, the performance of LLMs is highly sensitive to the prompt. Future research should explore the effects of different prompt configurations across the base models. Second, while this study focuses on ensemble of multiple models, it is necessary to examine the effect of combining different-scale models. This study treated all the outputs equally by using the median; however, to effectively integrate models with varying parameter counts, alternative ensemble techniques such as weighted averaging may be required. Third, although this study applied the ensemble of LLM inferences to marketing analysis on restaurant reviews, the proposed ensemble strategy itself is applicable to general text classification tasks. Future work should evaluate its effectiveness on broader benchmark across different domains to further validate the generalizability.